# Neural Solver Selection for Combinatorial Optimization

### v1

@@ 5 Conclusion and discussions
In this paper, we propose a general framework for neural solver selection for the first time, which can effectively select suitable solvers for each instance, leading to significantly better performance with little additional computational time, as validated by the extensive experiments on two well-studied COPs, TSP and CVRP. We hope this preliminary work can open a new line for the topic of neural combinatorial optimization. Within the proposed selection framework, we preliminarily investigate several implementations of the three key components: Feature extraction, training loss functions, and selection strategies. Techniques such as hierarchical graph encoder, ranking loss, rejection-based selection, and top-$p$ selection notably enhance overall performance. Beyond the techniques presented, we discuss several promising avenues for further research under this framework.
Feature extraction for neural solvers.
@@
In our implementation, we only extracted features for problem instances and used fixed indices for neural solvers, which assumes a static neural solver zoo and can not directly utilize any newly added neural solver during deployment. To enable zero-shot generalization to unseen neural solvers, it is essential to construct a smooth feature space for solvers, where those with similar preferences and biases are positioned closely together. Here we design a preliminary method for extracting features of neural solvers to facilitate generalization to unseen solvers. This method is based on the insight that a neural solver’s preferences can be characterized by representative instances where it significantly outperforms other solvers. For each neural solver, we sort those instances where it performs the best in an ascending order according to the ratio of the objective value that the solver obtains to the runner-up objective value, and select top 1% as its representative instances. Each representative instance is then treated as a token, and we apply a transformer to learn a summary feature from these instance tokens, which serves as the feature representation of the neural solver. Detailed implementations and results are provided in Appendix A.9. The results show that the preliminary method enables generalization to unseen neural solvers, where adding an extra solver can improve the selection performance.
Furthermore, advanced neural solver features should provide richer and deeper information than only using instance features to increase the capacity of the selection model. However, our preliminary method is based on the representative instances and fails to provide deeper information into the solvers’ internal mechanisms. For future improvements, some approaches may be worth exploring, such as utilizing large language models to encode the code of neural solvers (Wu et al., 2024) or learning neural representations from their trained parameters (Kofinas et al., 2024), which can access internal solver information, and potentially improving selection performance.
Runtime-aware selection for learn-to-seach solvers.
@@
In this paper, since the average runtime of most individual neural solvers is short (approximately 1–2 seconds), we ignored their time difference during the training of the selection model, and only used the objective values obtained by the neural solvers as supervision information (by classification or ranking). However, if there are some time-consuming learn-to-search solvers, such as NeuOpt (Ma et al., 2021; 2023) and local reconstruction methods (Kim et al., 2021; Ye et al., 2024), in the solver pool, the runtime should be considered in the performance ranking. In such cases, developing a runtime-aware selection method to balance computational time and solution optimality would be necessary. To address this, we could penalize objective values based on time consumption or simultaneously optimize both metrics using multi-objective learning methods (Lin et al., 2022).
Enhance the neural solver zoo by training.
@@
As shown in Figure 1, current neural solvers can exhibit complementary performance over instances without any modification, which has motivated our framework of neural solver selection. Inspired by the population-based training (Grinsztajn et al., 2023), we can further enhance their complementary ability through finetuning, i.e., each neural solver is finetuned on those instances where it performs the best. We can also train new solvers from scratch by maximizing their performance contribution to the current solver zoo and iteratively add such new solvers for enhancement. Moreover, to facilitate the training and deployment of a neural solver zoo, it is essential to develop a unified platform that provides interfaces for executing and training diverse neural solvers, such as an extension to the existing RL4CO (Berto et al., 2024).
@@ A.1 Graph attention layer
where $l$ is the layer index, $i$ is the node index, $\alpha^{l}$ is a learnable parameter used in the ReZero normalization, MHA and FF are short for the multi-head attention and the feed-forward network, respectively. For the implementations of the basic components MHA and FF, we refer to Vaswani et al. (2017) for details. Specifically, following the common settings of previous works, we set the dimension of $\bm{h}$ to 128, the number of heads in MHA to 8, and the hidden dimension of FF to 512.
A.2 Data generation
@@ A.4 Manual features
We reproduce the manual features proposed by Smith-Miles et al. (2010), which use statistical information and cluster analysis results to describe the characteristics of TSP. In this paper, we adopt these features: the standard deviation of the distances, the coordinates of the instance centroid, the radius of the TSP instance, the fraction of distinct distances, the variance of the normalized nearest neighbour distances (nNNd’s), the coefficient of variation of the nNNd’s, the ratio of the number of clusters to the number of nodes (Here we use HDBSCAN algorithm (Campello et al., 2013) to generate clusters), the ratio of number of outliers to nodes, and the mean radius of the clusters. For CVRP, we further add the mean and standard deviation of node demands to the features.
A.5 Results of top-$k$ selection.
@@
We compare the performance of our top-$k$ selection and the solver portfolio with the same size $k$ on four datasets, including TSP, CVRP, TSPLIB and CVRPLIB Set-X. As shown in Figure 3, our top-$k$ selection consistently outperforms the size-$k$ solver portfolio across $k\in\{1,2,3,4\}$. We also observe that the performance of our top-$k$ selection is close to the OPT when $k=4$. Moreover, it is expected that the performance improvement of our top-$k$ selection gradually diminishes as $k$ increases, since the performance of solver portfolio is also approaching the OPT (the gray line).
@@
A.6 Ablation of the hierarchical graph encoder.
The proposed hierarchical graph encoder utilizes graph pooling to downsample the instance graph and aggregates features obtained from multiple levels of the downsampled graphs. To evaluate the effectiveness of the graph pooling, we employ a graph encoder that aggregates the features from multiple layers for comparison, which is a clear ablation study since the main difference is that it does not have the graph pooling operation.
@@
A.7 Detailed results of rejection-based selection and top-$p$ selection.
The rejection-based selection and top-$p$ selection are both designed to achieve better performance with little additional time consumption. To evaluate them in detail, we tune their parameters (e.g., rejection ratio, $k$ and $p$) to obtain a range of results. For the rejection-based selection, we use $k\in\{2,3,4\}$ and vary the rejection ratio from 0.05 to 0.85 in increments of 0.05. For the top-$p$ selection, we adjust the the value of $p$ from 0.40 to 0.95 in increments of 0.01. The results of optimality gap and time consumption are provided in Figure 4. As shown in the figures, the rejection strategy with smaller $k$ tends to achieve better optimality gaps using the same time consumption. Therefore, we recommend $k=2$ or $3$ when using rejection-based selection. Comparing top-$p$ selection and rejection-based selection, we can not definitively conclude which strategy is superior, which is expected since they share similar idea of utilizing confidence levels to decide whether to employ multiple solvers. However, the top-$p$ selection may be preferable in practice due to its simplicity, where only a single hyperparameter $p$ requires tuning.
@@ A.8 Selection accuracy
A.9 Neural solver features
To enable generalization to unseen neural solvers, we design a preliminary feature extraction method for neural solver, which utilizes representative instances to represent a neural solver. For each neural solver, we sort those instances where the neural solver performs the best in an ascending order according to the ratio of the objective value that the solver obtains to the runner-up objective value, and select top 1% as its representative instances. Then, we use an instance encoder to obtain embeddings for each representative instance, serving as their token vectors. A transformer with two layers is employed to learn a summary feature from these instance tokens, which serves as the feature representation of the neural solver. More details are described as follows.
Instance tokenization.
We use a hierarchical graph encoder as the tokenization encoder to generate embeddings for each representative instance. Note that the parameters $\theta^{\prime}$ of the tokenization encoder are not updated by back-propagation. Instead, to stabilize the instance tokens during training, we update $\theta^{\prime}$ using a momentum-based moving average of the parameters $\theta$ of the instance feature encoder: $\theta^{\prime}\leftarrow m\cdot\theta^{\prime}+(1-m)\cdot\theta$, where $m\in[0,1)$ is a momentum coefficient (We set $m=0.99$ in experiments). Only the parameters $\theta$ are updated via back-propagation. This momentum update ensures that $\theta^{\prime}$ evolve more smoothly than $\theta$, resulting in stable instance tokenization.
Transformer architecture.
Selection model with neural solver feature.
The selection model integrates both the instance features and the neural solver features to output a score for each instance-solver pair. We employ an MLP to compute these scores. For each instance, the scores across all neural solvers are normalized to derive the probability distribution for solver selection.
To integrate a newly added neural solver, we first identify its representative instances from the training dataset and employ the aforementioned networks to compute its feature representation. The selection model can then leverage this new solver by considering its feature during selection, without the need for any fine-tuning.
To evaluate the effectiveness of this method, we remove the second best neural solver from the current solver zoo, train the selection model using the supervision information provided by the pruned solver zoo, and reintroduce the removed solver during testing. Figure 5 presents the top-$k$ selection performance with and without the newly added (extra) solver. The results show that the performance with newly added solver is generally better than the performance without it, demonstrating that the selection model can leverage the information of unseen solvers without any finetuning. In other words, the selection model can generalize to unseen solvers. However, we observe a slight decrease in top-$1$ performance, indicating that the preliminary method requires further improvement.
Figure 5: Performance of introducing an extra neural solver based on the neural solver feature. The results of top-$k$ selection with and without extra neural solver are presented.
A.10 Complete results with all neural solvers

### v2

@@ 5 Conclusion and discussions
In this paper, we propose a general framework for neural solver selection for the first time, which can effectively select suitable solvers for each instance, leading to significantly better performance with little additional computational time, as validated by the extensive experiments on two well-studied COPs, TSP and CVRP. Besides TSP and CVRP, our proposed selection framework is adaptable to other problems. For new problems, one only needs to customize the feature extraction component. For instance, when adapting our framework to scheduling problems, one can adjust the graph attention encoder according to MatNet (Kwon et al., 2021) (i.e., add edge embeddings). We hope this preliminary work can open a new line for neural combinatorial optimization. Within the proposed selection framework, we preliminarily investigate several implementations of the three key components: Feature extraction, training loss functions, and selection strategies. Techniques such as hierarchical graph encoder, ranking loss, rejection-based selection, and top-$p$ selection notably enhance overall performance. Beyond the techniques presented, we discuss several promising avenues for further research under this framework.
@@
Feature extraction for neural solvers. Our method, which uses fixed indices for neural solvers, assumes a static neural solver zoo and cannot directly utilize any newly added neural solvers during deployment. To enable zero-shot generalization to unseen neural solvers, it is essential to construct a smooth feature space for solvers, where those with similar preferences and biases are positioned closely together. In Section 4.3, we design a preliminary method for extracting features of neural solvers to facilitate generalization to unseen solvers. The results show that it enables generalization to unseen neural solvers, where adding an extra solver can improve the selection performance. For future improvements, some approaches may be worth exploring, such as utilizing large language models to encode the neural solvers from their codes and descriptions (Wu et al., 2024), or learning representations from their trained parameters (Kofinas et al., 2024), which can involve internal solver information and potentially bring improvements.
@@
Runtime-aware selection for learn-to-seach solvers. In this paper, since the average runtime of most individual neural solvers is short (approximately 1–2 seconds), we ignored their time difference during the training of the selection model, and only used the objective values obtained by the neural solvers as supervision information (by classification or ranking). However, if there are some time-consuming learn-to-search solvers, such as NeuOpt (Ma et al., 2021, 2023) and local reconstruction methods (Kim et al., 2021; Ye et al., 2024), in the solver pool, the runtime should be considered in the performance ranking. In such cases, developing a runtime-aware selection method to balance computational time and solution optimality would be necessary.
@@
Enhance the neural solver zoo by training. As shown in Figure 6, current neural solvers can exhibit complementary performance over instances without any modification, which has motivated our framework of neural solver selection. Inspired by the population-based training (Grinsztajn et al., 2023), we can further enhance their complementary ability through finetuning, i.e., each neural solver is finetuned on those instances where it performs the best. We can also train new solvers from scratch by maximizing their performance contribution to the current solver zoo and iteratively add such new solvers for enhancement. Moreover, to facilitate the training and deployment of a neural solver zoo, it is essential to develop a unified platform that provides interfaces for executing and training diverse neural solvers, such as an extension to the existing RL4CO (Berto et al., 2024).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgement
This work was supported by the Science and Technology Project of the State Grid Corporation of China (5700-202440332A-2-1-ZX).
@@ A.1 Graph attention layer
where $l$ is the layer index, $i$ is the node index, $\alpha^{l}$ is a learnable parameter used in the ReZero normalization, MHA and FF are short for the multi-head attention and the feed-forward network, respectively. For the implementations of the basic components MHA and FF, we refer to (Vaswani et al., 2017) for details. Specifically, following the common settings of previous works, we set the dimension of $\bm{h}$ to 128, the number of heads in MHA to 8, and the hidden dimension of FF to 512.
A.2 Supplement of experimental settings
A.2.1 Data generation
@@
A.2.2 Detailed Settings of open source solvers
We choose recent open-source neural solvers with state-of-the-art performance as the candidates, including Omni (Zhou et al., 2023), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), DIFUSCO (Sun & Yang, 2023), T2T (Li et al., 2023), ELG (Gao et al., 2024), INViT (Fang et al., 2024) and MVMoE (Zhou et al., 2024b). Greedy decoding is used for all the methods to avoid stochasticity. We set the pomo size to $100$ and the augmentation number to $8$ for the methods based on POMO (Kwon et al., 2020). The number of denoising steps is set to $50$ and the number of $2$-opt iterations is set to $100$ for diffusion-based methods. These individual solvers constitute a neural solver zoo. Ideally, if we can always select the best solver from the zoo for each instance, the optimal performance is achieved, which is also the performance upper bound of our selection model. Considering that some neural solvers contribute little to the overall performance, we iteratively eliminate the least contributive solver from the candidates, resulting in a more compact neural solver zoo, which reduces the zoo size to $7$ solvers for TSP and $5$ for CVRP. Further details of the elimination procedure are provided in Appendix A.3.
A.2.3 Detailed settings of hyperparameters
(1) Hyperparameters of graph encoders. For the graph attention encoder, we set the number of layers to $4$. For the hierarchical graph encoder, we use $2$ blocks where each block has $2$ attention layers. The embedding dimension is set to $128$. Other details of encoders can be found in Appendix A.1. (2) Hyperparameters of training. The Adam optimizer (Kingma & Ba, 2015) is employed for training, where we set the learning rate to $1\times 10^{-4}$ and the weight decay to $1\times 10^{-6}$. The number of epochs is set to $50$. The final model is chosen according to the performance on a validation dataset with $1,000$ synthetic instances. We train 5 selection models using different random seeds and report the mean and standard deviation of their performance. (3) Hyperparameters of selection strategies. For the top-$k$ strategy, we set $k=2$. For the rejection-based strategy, we reject the 20% of instances with the lowest confidence levels (i.e., the highest selection probability of all individual solvers), and apply top-$2$ selection to these rejected instances. For the top-$p$ strategy, we set $p=0.5$ for TSP and $p=0.8$ for CVRP.
@@ A.4 Manual features
We reproduce the manual features proposed by (Smith-Miles et al., 2010), which use statistical information and cluster analysis results to describe the characteristics of TSP. In this paper, we adopt these features: the standard deviation of the distances, the coordinates of the instance centroid, the radius of the TSP instance, the fraction of distinct distances, the variance of the normalized nearest neighbour distances (nNNd’s), the coefficient of variation of the nNNd’s, the ratio of the number of clusters to the number of nodes (Here we use HDBSCAN algorithm (Campello et al., 2013) to generate clusters), the ratio of number of outliers to nodes, and the mean radius of the clusters. For CVRP, we further add the mean and standard deviation of node demands to the features.
A.5 Comparisons with traditional algorithm selection methods
To further demonstrate the effectiveness of our proposed techniques, we provide additional comparison results between our proposed method and existing algorithm selection methods for non-neural TSP solvers (Smith-Miles et al., 2010; Seiler et al., 2020), as shown in Table 5. In fact, the method of using features from (Smith-Miles et al., 2010) and our ranking model was also compared in Table 3. The R package salesperson [^2] https//github.com/jakobbossek/salesperson provides the up-to-now most comprehensive collection of features for TSP and is widely used in algorithm selection methods (Seiler et al., 2020; Heins et al., 2021). Based on the feature set of salesperson, we reproduce an advanced algorithm selection method (Seiler et al., 2020) following the pipeline that computes hand-crafted features, conducts feature selection, and applies random forest for classification, where we employ the univariate statistical test to select important features. Besides, we also combine the salesperson features with our ranking model for ablation, denoted by ”(Seiler et al., 2020) + Ranking” in Table 5.
Table 5: Comparison experiments with algorithm selection methods for TSP. We report the mean (standard deviation) over five independent runs.
Methods Synthetic TSP TSPLIB
Gap Time Gap Time
Single best solver 2.33% 1.45s 1.95% 1.74s
Oracle 1.24% 8.93s 0.89% 9.14s
Algorithm selection methods
(Smith-Miles et al., 2010) + Ranking 1.97% (0.01%) 1.37s (0.01s) 1.83% (0.03%) 1.32s (0.05s)
(Seiler et al., 2020) 2.12% (0.04%) 1.35s (0.00s) 1.56% (0.01%) 1.34s (0.05s)
(Seiler et al., 2020) + Ranking 1.95% (0.01%) 1.33s (0.03s) 1.55% (0.03%) 1.27s (0.06s)
Our selection method using ranking
Greedy 1.86% (0.01%) 1.33s (0.01s) 1.33% (0.06%) 1.28s (0.03s)
Top-$k$ ($k=2$) 1.51% (0.02%) 2.56s (0.03s) 1.07% (0.03%) 2.48s (0.02s)
Rejection ($20\%$) 1.75% (0.02%) 1.63s (0.01s) 1.26% (0.03%) 1.51s (0.04s)
Top-$p$ ($p=0.5$) 1.68% (0.02%) 1.86s (0.07s) 1.28% (0.04%) 1.46s (0.06s)
The experimental results in Table 5 indicate that our proposed method can achieve superior performance than advanced algorithm selection methods on both synthetic TSP and TSPLIB. Comparing the fifth and sixth rows, our proposed hierarchical encoder demonstrates superior performance over the salesperson features, especially on the out-of-distribution benchmark TSPLIB. Additionally, the comparison of the fourth and fifth rows shows that our deep learning-based ranking model achieves better results than traditional classification methods. Furthermore, the results of the last three rows illustrate that our proposed adaptive selection strategies effectively enhance optimality with minimal increases in time consumption.
A.6 Results of top-$k$ selection.
@@ A.5 Comparisons with traditional algorithm selection methods
We compare the performance of our top-$k$ selection and the solver portfolio with the same size $k$ on four datasets, including TSP, CVRP, TSPLIB and CVRPLIB Set-X. As shown in Figure 4, our top-$k$ selection consistently outperforms the size-$k$ solver portfolio across $k\in\{1,2,3,4\}$. We also observe that the performance of our top-$k$ selection is close to the Oracle when $k=4$. Moreover, it is expected that the performance improvement of our top-$k$ selection gradually diminishes as $k$ increases, since the performance of solver portfolio is also approaching the Oracle (the gray line).
Related works, such as ZTop (Bai et al., 2021), employ a fixed set of neural solvers to construct a portfolio for all instances, resembling the static portfolio approach compared in this study. In contrast, our top-$k$ selection strategy dynamically constructs instance-specific portfolios, offering greater flexibility and a higher potential for performance improvement. As demonstrated in Figure 4, our method consistently outperforms the static portfolio approach across all portfolio sizes.
@@
A.7 Ablation of the hierarchical graph encoder.
The proposed hierarchical graph encoder utilizes graph pooling to downsample the instance graph and aggregates features obtained from multiple levels of the downsampled graphs. A graphical illustration of the downsampling process is provided in Appendix A.11, where we can find some consistent patterns which are intuitively reasonable. To evaluate the effectiveness of the graph pooling, we employ a graph encoder that aggregates the features from multiple layers for comparison, which is a clear ablation study since the main difference is that it does not have the graph pooling operation.
@@
To evaluate the computational efficiency of the hierarchical encoder, we provide detailed comparisons of the computation cost and optimality between our hierarchical encoder and a typical graph encoder. The results are shown in Table 7, which includes the inference time per instance on TSPLIB, training time per epoch, and the average optimality gap on TSPLIB.
Table 7: Comparisons of the computation cost and optimality between our hierarchical encoder and a typical attention encoder.
Methods Inference time of selection model Inference time of neural solvers Training time each epoch Optimality gap
Naive attention encoder 0.0054s 1.2600s 1m40s 1.54%
Hierarchical encoder 0.0070s 1.2961s 2m30s 1.37%
We can observe from the second column that the introduction of our hierarchical encoder will increase the inference time of the selection model a little bit, e.g., from 0.0054s to 0.0070s. However, as shown in the second and third columns, the inference time of the selection model is orders of magnitude shorter than that of the neural solvers, so the inference efficiency of the selection model is less of a concern. The fourth column shows that the training time per epoch of the naïve encoder and the hierarchical encoder are 1m40s and 2m30s, respectively. Although the hierarchical encoder slows the training, the total runtime for 50 epochs is still only 2 hours, which is acceptable in most scenarios. Therefore, the performance metric (i.e., optimality gap) of different encoders is more crucial, especially the generalization performance. If the encoder learns robust representations, we can directly transfer the selection model to different datasets in a zero-shot manner, saving the time for fine-tuning and adaptation. Considering the better generalization (e.g., the optimality gap decreases from 1.54% to 1.37%), we believe that the proposed hierarchical encoder is a better choice.
A.8 Detailed comparisons of selection strategies
According to the mechanisms of the four selection strategies, they have different preferences in the trade-off of efficiency and optimality. Generally, for efficiency, Greedy $>$ Rejection $\approx$ Top-$p$ $>$ Top-$k$, for optimality, Top-$k$ $>$ Rejection $\approx$ Top-$p$ $>$ Greedy. Meanwhile, the hyper-parameters of them can be used for balancing efficiency and optimality as well. As a result, the choice of different selection strategies can be decided by the users according to their preference, and we suggest using Top-p or Rejection as the default choices since they can adaptively select solvers based on the confidence.
The rejection-based selection and top-$p$ selection are both designed to achieve better performance with little additional time consumption. To evaluate them in detail, we tune their parameters (e.g., rejection ratio, $k$, and $p$) to obtain a range of results. For the rejection-based selection, we use $k\in\{2,3,4\}$ and vary the rejection ratio from 0.05 to 0.85 in increments of 0.05. For the top-$p$ selection, we adjust the value of $p$ from 0.40 to 0.95 in increments of 0.01. The results of the optimality gap and time consumption are provided in Figure 5. As shown in the figures, the rejection strategy with smaller $k$ tends to achieve better optimality gaps using the same time consumption. Therefore, we recommend $k=2$ or $3$ when using rejection-based selection. Comparing top-$p$ selection and rejection-based selection, we can not definitively conclude which strategy is superior, which is expected since they share a similar idea of utilizing confidence levels to decide whether to employ multiple solvers. However, the top-$p$ selection may be preferable in practice due to its simplicity, where only a single hyperparameter $p$ requires tuning.
@@ A.9 Selection accuracy
A.10 Additional results for more neural solvers and larger-scale datasets
We add two divide-and-conquer solvers, GLOP (Ye et al., 2024) and UDC (Zheng et al., 2024), to our solver pool, increase the problem scale from $N\in[50,500]$ to $N\in[500,2000]$ to conduct new experiments. The results shown in Table 9 demonstrate that our framework can be compatible with more neural solvers and can also improve performance over the single best solver on larger-scale instances.
Table 9: Experimental results on the larger-scale instances with $N\in[500,2000]$. We report the mean (standard deviation) over five independent runs.
Methods Synthetic TSP with $N\in[500,2000]$
Gap Time
Single best solver 6.104% 8.369s
Ours (Greedy) 5.540% (0.038%) 8.322s (0.036s)
Ours (Top-$k$, $k=2$) 5.369% (0.003%) 15.566s (0.085s)
Single best of new solver pool 3.562% 5.274s
Ours with new solvers (Greedy) 3.126% (0.002%) 6.892s (0.006s)
Ours with new solvers (Top-$k$, $k=2$) 2.955% (0.005%) 13.713s (0.036s)
A.11 Illustration of the nodes sampled by hierarchical graph encoder
We illustrate the retained nodes after downsampling. Surprisingly, we can find some consistent patterns which are intuitively reasonable. We summarize them as three main points:
  • Cluster nodes. As illustrated in Figures 6 and 6, when instances contain certain clusters, the hierarchical encoder tends to select a subset of “representative” nodes from each cluster, efficiently describing the entire spatial distribution.
  • Specific blocks. As illustrated in Figures 6 and 6, when instances contain specific complex geometric patterns like squares (Figure 6) and arrays (Figure 6), the hierarchical encoder can capture the nodes of these important areas to identify their characteristics.
  • Boundary nodes. For instances without clear sub-components, the hierarchical encoder tends to focus on boundary nodes that describe the global shape, as illustrated in Figures 6 and 6.
Figure 6: Illustrations of nodes selected by the hierarchical encoder. Each sub-figure represents an instance of TSP. The blue nodes represent the original instance, and the red nodes represent the retained nodes after down-sampling by the hierarchical encoder.
A.12 Complete results with all neural solvers
The complete experimental results with all individual solvers are presented in Table 10 and Table 11, where the average objective values are provided for further comparisons.